# The canonical lamination calibrated by a cohomology class

### v3

@@ 5.5. Strict convexity and the derived series of $\pi_{1}(M)$
The next two lemmata appeared in [Auer12], though the proof of Theorem 1.6(2) does not. Since this manuscript is not publicly available, or complete, we reproduce them here with full credit to the original authors.
Lemma 5.9 ([Auer12]).
@@
Let $F$ be a fundamental domain of $M$ in $\tilde{M}^{\rm ab}$. Since $u^{\rm ab}\in L^{\infty}_{\mathrm{loc}}$ [Gorny20, Theorem 4.3] and $F$ is compact, there exists $t_{0}\in\mathbf{R}$ such that $u>t_{0}$ on $F$. Using (5.2) to interpret $H_{1}(M,\mathbf{Z})$ as the group of deck transformations of $\tilde{M}^{\rm ab}$, let
@@
By a straightforward generalization of [BOMBIERI1969, Theorem 1], $\hat{X}$ is perimeter-minimizing. If $\hat{X}$ is bounded, then $\partial\hat{X}$ is competing with the empty set and hence is empty, a contradiction; so $\hat{X}$ is unbounded and therefore has infinite volume by Lemma 5.8. But $\psi$ is an isometry, so $\psi_{*}(\hat{X})$ is an infinite-volume subset of the closed manifold $M$, a contradiction. ∎
Lemma 5.10 ([Auer12]).
@@
We reason by contrapositive. Let $N_{1},N_{2}$ be two distinct components of $\partial\{u^{\rm ab}>t\}$. By Lemma 5.9 (and the analogous result for sublevel sets), $\tilde{M}^{\rm ab}\setminus\partial\{u^{\rm ab}>t\}$ has two components $E_{1},E_{2}$. We construct a curve $\gamma$, transverse to $N_{1}$, which starts at a point $x\in N_{1}$, passes through $E_{1}$, crosses $N_{2}$ into $E_{2}$, and then returns to $x$. In particular $\gamma$ meets $N_{1}$ at a single point, so their intersection number $[\gamma]\cdot[N_{1}]=1$ (possibly after reorienting). Therefore $[\gamma]$ is a nontrivial class in $H_{1}(\tilde{M}^{\rm ab},\mathbf{R})$. ∎
@@ 5.6. The earthquake norm
The picture which seems to be emerging is that the stable norm is highly analogous to the earthquake norm on the cotangent space $T_{\sigma}^{*}\mathscr{T}_{g}$ to the Teichmüller space of a closed hyperbolic surface $(\Sigma_{g},\sigma)$ of genus $g$. The starting point for this observation is the earthquake theorem, [Kerckhoff83], which asserts that one can identify each $\alpha\in T_{\sigma}^{*}\mathscr{T}_{g}$ uniquely with a measured geodesic lamination in $(\Sigma_{g},\sigma)$; the earthquake norm $\|\alpha\|_{\rm eq}$ is the mass of the corresponding measure. It follows from [daskalopoulos2023, Theorem 1.6] that for each geodesic lamination $\lambda$ in $(\Sigma,\sigma)$, the map $\mathcal{M}(\lambda)\to T_{\sigma}^{*}\mathscr{T}_{g}$ is affine. The dual norm to the earthquake norm is the stretch norm, which is given by infinitesimal minimizing Lipschitz maps, just as the costable norm is given by calibrations.
@@
Theorem 5.11 ([huang2024earthquakemetricteichmullerspace, Theorem 6.1]).
@@
The intersection product on measured geodesic laminations on $(\Sigma_{g},\sigma)$ corresponds to the Weil-Petersson $2$-form; in particular, it is symplectic. So by Theorem 5.11, every earthquake flat $F$ is contained in a $33g-$ 3-dimensional subspace of $T_{\sigma}^{*}\mathscr{T}_{g}$. Therefore the earthquake norm also satisfies an analogue of Theorem 1.4:
@@
    $F$ is a convex polytope with at most $33g-$ 3 vertices.
@@
Let $(\alpha_{i})$ be a dense sequence in $F$. By Theorem 5.11, if we let $\lambda_{i}:=\bigcup_{j\leq i}\operatorname{supp}\alpha_{j}$, then $\lambda_{i}$ is (the support of) a geodesic lamination such that $\lambda_{i}\subseteq\lambda_{i+1}$. Taking the limit (say, in the Vietoris topology), $\lambda_{i}$ converges to a lamination $\lambda$ for which $\alpha_{i}\in\mathcal{M}(\lambda)$. Since $\mathcal{M}(\lambda)$ is compact, $F\subseteq\mathcal{M}(\lambda)$, and since $F$ is maximal, $\mathcal{M}(\lambda)\subseteq F$. By an easy generalization of Theorem 5.5, $\mathcal{E}(F)$ is the set of ergodic measures on $\mathcal{M}(\lambda)$, so it is linearly independent. But $\mathcal{E}(F)$ is contained in a $33g-$ 3-dimensional vector space, so $\operatorname{card}\mathcal{E}(F)\leq 3g-3$. ∎
@@
Conjecture 5.13 seems quite likely to hold in view of Corollaries 1.3 and 5.12. A natural attempt to prove Conjecture 5.13 is to show that there is a diffeomorphism $\exp_{\sigma}:T_{\sigma}\mathscr{T}_{g}\to\mathscr{T}_{g}$ which maps every ray emanating from 00 in $T_{\sigma}\mathscr{T}_{g}$ to a geodesic ray in $\mathscr{T}_{g}$ with respect to Thurston’s stretch metric, such that $v^{*}\subseteq\mathcal{M}(\lambda_{\sigma,\exp_{\sigma}(v)})$. We refer to Pan and Wolf, [pan2022raystructuresteichmullerspace], for more discussion of exponential maps for Thurston’s stretch metric.
[…]

### v4

@@ 5.4. Strict convexity and the derived series of $\pi_{1}(M)$
The next two lemmata appeared in [4], though the proof of Theorem 1.6(2) does not. Since this manuscript is not publicly available, or complete, we reproduce them here with full credit to the original authors.
Lemma 5.8 ([4]).
@@
Let $F$ be a fundamental domain of $M$ in $\tilde{M}^{\rm ab}$. Since $u^{\rm ab}\in L^{\infty}_{\mathrm{loc}}$ [26, Theorem 4.3] and $F$ is compact, there exists $t_{0}\in\mathbf{R}$ such that $u>t_{0}$ on $F$. Using (5.1) to interpret $H_{1}(M,\mathbf{Z})$ as the group of deck transformations of $\tilde{M}^{\rm ab}$, let
@@
[…]
Lemma 5.9 ([4]).
@@
We reason by contrapositive. Let $N_{1},N_{2}$ be two distinct components of $\partial\{u^{\rm ab}>t\}$. By Lemma 5.8 (and the analogous result for sublevel sets), $\tilde{M}^{\rm ab}\setminus\partial\{u^{\rm ab}>t\}$ has two components $E_{1},E_{2}$. We construct a curve $\gamma$, transverse to $N_{1}$, which starts at a point $x\in N_{1}$, passes through $E_{1}$, crosses $N_{2}$ into $E_{2}$, and then returns to $x$. In particular $\gamma$ meets $N_{1}$ at a single point, so their intersection number $\gamma]\cdot[N_{1}]=1$ (possibly after reorienting). Therefore $\gamma]$ is a nontrivial class in $H_{1}(\tilde{M}^{\rm ab},\mathbf{R})$. ∎
@@ 6.1. The earthquake norm
The picture which seems to be emerging is that the stable norm is highly analogous to the earthquake norm on the cotangent space $T_{\sigma}^{*}\mathscr{T}_{g}$ to the Teichmüller space of a closed hyperbolic surface $(\Sigma_{g},\sigma)$ of genus $g$. The starting point for this observation is the earthquake theorem, [33], which asserts that one can identify each $\alpha\in T_{\sigma}^{*}\mathscr{T}_{g}$ uniquely with a measured geodesic lamination in $(\Sigma_{g},\sigma)$; the earthquake norm $\|\alpha\|_{\rm eq}$ is the mass of the corresponding measure. It follows from [16, Theorem 1.6] that for each geodesic lamination $\lambda$ in $(\Sigma,\sigma)$, the map $\mathcal{M}(\lambda)\to T_{\sigma}^{*}\mathscr{T}_{g}$ is affine. The dual norm to the earthquake norm is the stretch norm, which is given by infinitesimal minimizing Lipschitz maps, just as the costable norm is given by calibrations.
@@
Theorem 6.1 ([31, Theorem 6.1]).
@@
The intersection product on measured geodesic laminations on $(\Sigma_{g},\sigma)$ corresponds to the Weil-Petersson $2$-form; in particular, it is symplectic. So by Theorem 6.1, every earthquake flat $F$ is contained in a 3g33g-3-dimensional subspace of $T_{\sigma}^{*}\mathscr{T}_{g}$. Therefore the earthquake norm also satisfies an analogue of Theorem 1.4:
@@
    $F$ is a convex polytope with at most 3g33g-3 vertices.
@@
Let $(\alpha_{i})$ be a dense sequence in $F$. By Theorem 6.1, if we let $\lambda_{i}:=\bigcup_{j\leq i}\operatorname{supp}\alpha_{j}$, then $\lambda_{i}$ is (the support of) a geodesic lamination such that $\lambda_{i}\subseteq\lambda_{i+1}$. Taking the limit (say, in the Vietoris topology), $\lambda_{i}$ converges to a lamination $\lambda$ for which $\alpha_{i}\in\mathcal{M}(\lambda)$. Since $\mathcal{M}(\lambda)$ is compact, $F\subseteq\mathcal{M}(\lambda)$, and since $F$ is maximal, $\mathcal{M}(\lambda)\subseteq F$. By an easy generalization of Theorem 5.5, $\mathcal{E}(F)$ is the set of ergodic measures on $\mathcal{M}(\lambda)$, so it is linearly independent. But $\mathcal{E}(F)$ is contained in a 3g33g-3-dimensional vector space, so $\operatorname{card}\mathcal{E}(F)\leq 3g-3$. ∎
@@
Conjecture 6.3 seems quite likely to hold in view of Corollaries 1.3 and 6.2. A natural attempt to prove Conjecture 6.3 is to show that there is a diffeomorphism $\exp_{\sigma}:T_{\sigma}\mathscr{T}_{g}\to\mathscr{T}_{g}$ which maps every ray emanating from 0 in $T_{\sigma}\mathscr{T}_{g}$ to a geodesic ray in $\mathscr{T}_{g}$ with respect to Thurston’s stretch metric, such that $v^{*}\subseteq\mathcal{M}(\lambda_{\sigma,\exp_{\sigma}(v)})$. We refer to Pan and Wolf, [41], for more discussion of exponential maps for Thurston’s stretch metric.
6.2. Higher dimension and codimension
In this paper we only considered cohomology classes in $H^{d-1}(M,\mathbf{R})$ where the manifold $M$ had dimension $d\leq 7$. However, the definition of the costable norm makes sense in any cohomology group $H^{k}(M,\mathbf{R})$ and so it is natural to ask if Theorem 1.2 holds for $k\leq d-2$ or $d\geq 8$.
A version of Theorem 1.2 is available when $k=1$. Let $\rho\in H^{1}(M,\mathbf{R})$ and $\|\rho\|_{\infty}=1$. By thinking of $\rho$ as a representation $\pi_{1}(M)\to\mathbf{R}$, we can identify calibrations in $\rho$ with functions $v$ on $\tilde{M}$ such that $\mathrm{Lip}(v)=1$. From this, and the proof of [28, Lemma 5.2], it follows that there is a unique largest geodesic lamination $\lambda_{\rho}$ which is calibrated by every calibration in $\rho$. It seems plausible that one can generalize arguments of [17, §6], to establish that $\rho^{*}$ is the set of homology classes of transverse measures to sublaminations of $\lambda_{\rho}$, and show that the rational extreme points of $\rho^{*}$ correspond to closed leaves in $\lambda_{\rho}$. However, if $d\geq 3$ then the codimension of $\lambda_{\rho}$ is too large to hope that any of the intersection-theoretic results of this paper generalize to the $k=1$ case.
If $2\leq k\leq d-2$, then there can be no generalization of Theorem 1.2 to $H^{k}(M,\mathbf{R})$, since the submanifolds calibrated by a class $\rho\in H^{k}(M,\mathbf{R})$ can intersect each other. For example, let $M=\mathbf{P}^{2}_{\mathbf{C}}$, let $\rho$ be the Kähler class of $\mathbf{P}^{2}_{\mathbf{C}}$, and let $F$ be the Kähler form of $\mathbf{P}^{2}_{\mathbf{C}}$, which is a calibration by Wirtinger’s inequality. Then every projective algebraic curve in $\mathbf{P}^{2}_{\mathbf{C}}$ is $F$-calibrated, and so is a leaf of the putative canonical lamination calibrated by $\rho$!
The construction of the canonical calibrated lamination heavily relied on the regularity theory of stable minimal hypersurfaces [44], which is only valid when $d\leq 7$. We outline a possible approach to generalizing this paper to $d\geq 8$. Let $F$ be a calibration, $\rho:=[F]$, $N$ be an $F$-calibrated hypersurface, and $x\in N$. After throwing away a set of codimension $\geq 8$, we may assume that $x$ is not a singular point of $N$, so that $|\mathrm{I\!I}_{N}|$ is bounded near $x$. Let $N^{\prime}$ be another $F$-calibrated hypersurface such that $x$ is not a singular point of $N^{\prime}$. It should follow from Simon’s maximum principle [45] and the proof of Lemma 4.4 that there exists $\delta>0$ such that $B(x,\delta)\cap N\cap N^{\prime}=\emptyset$.
The natural next step in the construction of $\lambda_{\rho}$ would be to assume towards contradiction that $N^{\prime}$ oscillates wildly at some point $y\in N^{\prime}\cap B(x,\delta)$. By elliptic regularity, and the fact that $N,N^{\prime}$ are disjoint, one would like to conclude that $N$ also oscillates wildly, contradicting the bound on $|\mathrm{I\!I}_{N}|$. If that argument worked, then we would conclude that $|\mathrm{I\!I}_{N^{\prime}}|$ is bounded near $y$. By Allard’s regularity theorem [12, Theorem 7.8], that argument almost works, but actually only shows that if $|\mathrm{I\!I}_{N^{\prime}}|$ is very large at some point close to $y$, then on coarse scales, $N^{\prime}$ looks like a current of higher multiplicity at $y$. This is the same mechanism by which a stack of catenoids (Example 2.9) fails to be a lamination.
It is possible that stacks of catenoids constitute the worst-case scenario. Motivated by this, we make two bold conjectures.
Conjecture 6.4.
Let $\mathcal{S}$ be a set of disjoint minimal hypersurfaces in $M$. Then there is a closed set $Z$ such that $\dim Z\leq d-3$, and for every $U\Subset M$ there is a Lipschitz lamination $\lambda$ in $U$ such that $\{N\cap U:N\in\mathcal{S}\}$ is dense in the set of leaves of $\lambda$.
Conjecture 6.5.
Let $d\geq 8$, let $\rho\in H^{d-1}(M,\mathbf{R})$, assume that $\|\rho\|_{\infty}=1$, and let $\mathcal{S}$ be the set of minimal hypersurfaces which are calibrated by every calibration in $\rho$. Then there is a closed set $Z$ such that $\dim Z\leq d-3$ and, for every $U\Subset M\setminus Z$, there is a Lipschitz lamination $\lambda$ in $U$ such that $\{N\cap U:N\in\mathcal{S}\}$ is dense in the set of leaves of $\lambda$.
The results of §5 should not be able to detect the presence of a codimension $\geq 3$ singular set, so we expect them to generalize ceteris paribus to $d\geq 8$.